OpenStudy (anonymous):

*help requested* *fan and medal awarded* Solve for x: -5(x - 3) = 30 x = 9 x = -9 x = 3 x = -3 4x < -16 or x + 5 >= 15 A number line with a closed circle on -4, shading to the left, and an open circle on 10, shading to the right. A number line with an open circle on -4, a closed circle on 10, and shading in between. A number line with a closed circle on -4, an open circle on 10, and shading in between. A number line with an open circle on -4, shading to the left, and a closed circle on 10, shading to the right.

OpenStudy (anonymous):

its x= 8 ,9=4 x=4

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

-5x + 15 = 30 15 15 -5x =15 -5 -5 x = -3

OpenStudy (anonymous):

thats the first one

OpenStudy (anonymous):

thanks how bout the other one ?

OpenStudy (anonymous):

btw i forget open circle is < and > while closed circle is <= and >=

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

srry i meant did u know :P

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

Part 1: 4x < -16 /4 /4 x < -4 Part 2: x + 5 >= 15 -5 -5 x >= 10 so your answer is... D. A number line with an open circle on -4, shading to the left, and a closed circle on 10, shading to the right.

OpenStudy (anonymous):

can u help with another ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which ordered pair is a solution to the inequality y > x + 5? (-3, 2) (0, -3) (3, 5) (2, 8)

OpenStudy (anonymous):

Let A = {t-shirt, tank, oxford} and let B = {shorts, pants, skirt}. Find A X B. {t-shirt, tank, oxford, shorts, pants, skirt} { } {(t-shirt, shorts), (t-shirt, pants), (t-shirt, skirt), (tank, shorts), (tank, pants), (tank, skirt), (oxford, shorts), (oxford, pants), (oxford, skirt)} {(t-shirt, short), (tank, pants), (oxford, skirt)}

OpenStudy (anonymous):

i can do the second one the other im not sure of

OpenStudy (anonymous):

c